OpenStudy (anonymous):

given z=√2-√2 i how to solve z^13. i have no clue where to start

OpenStudy (kirbykirby):

Express the complex number in polar form. If \(z=a+bi\), then \(\large z=re^{i\theta}\) where \(r=\sqrt{a^2+b^2}\) and \(\tan\theta =\dfrac{b}{a}\) And then use DeMoivre's theorem, \(z^n =(re^{i\theta})^n=r^n(\cos(n\theta)+i\sin (n\theta))\)

OpenStudy (anonymous):

could you perhaps show an example with the given numbers?, im not sure i understand it at all

OpenStudy (kirbykirby):

\(z=\sqrt{2}-i\sqrt{2}\) and \(z=re^{i\theta}\) \(r=\sqrt{\left(\sqrt{2}\right)^2+\left(-\sqrt{2}\right)^2}=\sqrt{2+2}=2\) \(\tan\theta=\dfrac{-\sqrt{2}}{\sqrt{2}}=-1\implies\theta= \arctan(-1)=\dfrac{-\pi}{4}\) So... \(\large z=re^{i\theta}=2e^{-i\frac{\pi}{4}}\) Recalling that \(e^{i\theta}=cos\theta+i\sin\theta\)... then \(\large 2e^{-i\frac{\pi}{4}}=2\left[\cos\left(\frac{-\pi}{4}\right)+i\sin\left(\frac{-\pi}{4}\right)\right]\) By DeMoivre's theorem: \(z^n =(re^{i\theta})^n=r^n(\cos(n\theta)+i\sin (n\theta))\), so: \(\large \left(2e^{-i\frac{\pi}{4}}\right)^{13}=2^{13}\left[\cos\left(\frac{-13\pi}{4}\right)+i\sin\left(\frac{-13\pi}{4}\right)\right]\) Now, you can convert the \(\dfrac{-13\pi}{4}\) into \(\dfrac{3\pi}{4}\) by adding 2 multiples of \(2\pi\) Cn you proceed from here?

OpenStudy (anonymous):

hmm im not sure, i already feel bad for nagging..but if you wouldnt mind it would be nice to have something to compare whatever answer i may get to.

OpenStudy (anonymous):

im sitting on (-5792.62 + 5792,62i) as an answer

OpenStudy (kirbykirby):

yes :), although I think it's preferable to use \(\sqrt{2}\) in your answer instead of putting it in decimal form ;)

OpenStudy (anonymous):

hmm noted! thanks a lot for your time and effort in teaching me this!=) you saved my assignment;p

OpenStudy (anonymous):

btw how do you come to that answer? i tried and it became a mess;p

OpenStudy (anonymous):

the √2 answer that is

OpenStudy (kirbykirby):

hm I suspect you just plugged that in your calculator lol.

OpenStudy (kirbykirby):

Recall from the unit circle that \[\cos\left(\frac{\pi}{4}\right)=\frac{\sqrt{2}}{2} \] and \[\sin\left(\frac{\pi}{4}\right)=\frac{\sqrt{2}}{2} \] So: To find \[ \cos\left(\frac{3\pi}{4}\right)\] and \[ \cos\left(\frac{3\pi}{4}\right)\], use symmetry of the unit circle: